is considering adding state safety inspections to its service offerings. The equipment necessary to perform these inspections will cost $512,000 and will generate cash flows of $179,000 over each of the next five years. If the cost of capital is 18 percent, what is the MIRR on this project?
Answer #1

MIRR = ( FVc / PVfc )1/n -1

where, FVc is the  future value of positive cash flows

and PVfc is the present value of negative cash flows

CF0 = -512000
CF1 = 179000
CF2 = 179000
CF3 = 179000
CF4 = 179000
CF5 = 179000

Cost of capital = r = 18%

PVfc = 512000

FVc = ΣCFn(1+r)t = 179000*1.184 + 179000*1.183 + 179000*1.182 + 179000*1.18 + 179000 = 1280603.55

n is the cash flow in year n and t is the years to maturity

Hence, MIRR = (1280603.55/512000)1/5 - 1 = 0.2012 or 20.12%